OpenStudy (anonymous):

If 10^k = 64, what is the value of 10^(k/2 +1)?

OpenStudy (anonymous):

18 42 80 81 320

OpenStudy (kirbykirby):

Ok so it's a matter of breaking it down:

OpenStudy (kirbykirby):

You need to be familiar with the laws of exponents. \[10^{\frac{k}{2}+1}\]... do you know the law \(a^b*a^c=a^{b+c}\)

OpenStudy (anonymous):

yes

OpenStudy (kirbykirby):

so here it's the same idea, but here a=10 and "k/2"=b and "1"=c

OpenStudy (kirbykirby):

can you re-write it using the law?

OpenStudy (anonymous):

10^(k/2)+ 10^1

OpenStudy (kirbykirby):

almost... just replace your addition + with a multiplication

OpenStudy (anonymous):

10^(k/2) x 10^1

OpenStudy (kirbykirby):

yes, now there is another law \[(a^B)^C=a^{BC}\] and we can apply it to \[10^{\frac{k}{2}}\] by noticing that \[\frac{k}{2}=k*\frac{1}{2}\]

OpenStudy (anonymous):

so 320?

OpenStudy (kirbykirby):

Nope =\ If you re-write \[10^{\frac{k}{2}}=10^{k*\frac{1}{2}}\]

OpenStudy (kirbykirby):

You can see that you can say \((10^k)^{\frac{1}{2}}\)

OpenStudy (anonymous):

First we notice the following:\[\bf 10^{\frac{k}{2}+1}=10^{\frac{k}{2}}(10)=(10^k)^\frac{1}{2}(10)=10\sqrt{10^k}\]We know that \(\bf 10^k=64\), so plug that in and evaluate.

OpenStudy (anonymous):

@sakigirl